OpenStudy (anonymous):

Ok here is another Simplify question I do need it broken down cause its not making sense. Cube square root of -27

OpenStudy (lgbasallote):

your question is \[\Large \sqrt[3]{-27}\] right? here's a hint \[1^3 = 1\] \[2^3 = 8\] \[3^3 =27\] \[4^3 = 64\] \[5^3 = 125\] etc. does that help?

OpenStudy (anonymous):

she is asking about the negative part i think it will be following what is called principal root so all answers is positive right??? So its 3 I guess

OpenStudy (hba):

3i (i=iota)

OpenStudy (anonymous):

iota??? what does that mean

OpenStudy (anonymous):

\[\sqrt[3]-27 =\sqrt[3]{-9 } x 3 = \]

OpenStudy (lgbasallote):

@mandypruitt you're pretty close.. \[\LARGE \sqrt[3]{-27} \implies \sqrt[3]{-3 \times -3 \times -3} \implies \sqrt[3]{(-3)^3}\] can you see it now?

OpenStudy (anonymous):

\[\sqrt[3]{-9} . \sqrt[3]{3}\]

OpenStudy (lgbasallote):

hint: \[\huge \sqrt[m]{(a)^m} \implies a\]

OpenStudy (anonymous):

try to recall the rules about the sign when multiplying and dividing negative and positive integers xD

OpenStudy (anonymous):

I forsome reason cant get passed \[\sqrt[3]{-9} \times \sqrt[3]{3}\]

OpenStudy (lgbasallote):

@mandypruitt do what i wrote above

OpenStudy (lgbasallote):

i assure you that where you are right now is a dead end

OpenStudy (anonymous):

Im sorry but I am at a loss

OpenStudy (lgbasallote):

okay..let's start from the beginning..

OpenStudy (lgbasallote):

do you agree that 27 = 9 x 3?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

what about 27 = 3 x 3 x 3 do you agree that is right?

OpenStudy (anonymous):

yes

OpenStudy (jiteshmeghwal9):

U can write it as \[\LARGE{\sqrt[3]{-1 \times 27}}\]\[\LARGE{\sqrt[3]{-1} \times \sqrt[3]{27}}\]\[\LARGE{3i}\]

OpenStudy (lgbasallote):

so -27 = -3 x -3 x -3 right? -3 x - 3 is +9 then 9 x - 3 is -27 following so far?

OpenStudy (anonymous):

yes

OpenStudy (jiteshmeghwal9):

now \(\LARGE{3i=3 \times -1=-3 Ans.}\)

OpenStudy (anonymous):

-3 * -3 * -3 = ???

OpenStudy (lgbasallote):

so \[\Large \sqrt[3]{-27} = \sqrt[3]{-3 \times -3 \times -3}\] right?

OpenStudy (jiteshmeghwal9):

or simply do it as \[-3 \times -3 \times -3\]

OpenStudy (jiteshmeghwal9):

take common of them

OpenStudy (jiteshmeghwal9):

-3 ans

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

so now.. according to algebra \[-3 \times -3 \times - 3 = (-3)^3\] right?

OpenStudy (anonymous):

yes

OpenStudy (hba):

no the answer is 3i^2

OpenStudy (hba):

i^2 =-1 where i=iota

OpenStudy (anonymous):

-3*-3*-3 =-27 (-)(-)(-) = (+)(-) = (-)

OpenStudy (lgbasallote):

so \[\Large \sqrt[3]{-3 \times -3 \times -3} \implies \sqrt[3]{(-3)^3}\] you're still following right?

OpenStudy (anonymous):

ok cuberoot of -27 = -3 xD

OpenStudy (anonymous):

yes

OpenStudy (hba):

I am right the answer is 3i^2

OpenStudy (lgbasallote):

good.. now refer to the algebra rule i posted a while ago \[\huge \sqrt[m]{a^m} \implies a\] so what do you think is the value of \[\huge \sqrt[3]{(-3)^3}\]

OpenStudy (anonymous):

-3 ONLY

OpenStudy (hba):

It Becomes Complex guys

OpenStudy (jiteshmeghwal9):

\[\LARGE\sqrt{(-3)^3}{}\]\[\LARGE{(-3)^{3\times 1\over 3}}\]

OpenStudy (jiteshmeghwal9):

-3

OpenStudy (hba):

no ^ 3i^2 is the answer i can challenge u guys

OpenStudy (lgbasallote):

maybe you guys should debate somewhere else?

OpenStudy (hba):

My answer is correct dude

OpenStudy (anonymous):

no its not and leave convo

OpenStudy (jiteshmeghwal9):

hey guys -3 is the answer \(\LARGE100 percent\)

OpenStudy (hba):

Have you guys studied complex numbers

OpenStudy (jiteshmeghwal9):

yes!

OpenStudy (lgbasallote):

please debate somewhere else =_= it's making the post laggy

OpenStudy (hba):

And You dont know what to answer when u see see a negative sign in underroot

OpenStudy (anonymous):

you guys are just confusing me more so I guess I will figure it out somewhere else.

OpenStudy (hba):

I am out sorry for worrying you @mandypruitt

OpenStudy (jiteshmeghwal9):

we should write it as \(i\)

OpenStudy (jiteshmeghwal9):

@hba :)

OpenStudy (lgbasallote):

to tell you frankly...they're debating on the same answer :|

OpenStudy (hba):

^^ Thats what i am saying

OpenStudy (anonymous):

I agree its 3i

OpenStudy (lgbasallote):

\[3i^2 = 3(-1) = -3\] =_= why the heck are you guys fighting over the same answer

OpenStudy (jiteshmeghwal9):

\[\LARGE{\sqrt[3]{-1 \times 27}}\]\[\LARGE{\sqrt[3]{-1} \times \sqrt[3]{27}}\]

OpenStudy (lgbasallote):

anyway @mandypruitt \[\large \sqrt[3]{-27} \implies \sqrt[3]{-3 \times -3 \times -3} \implies \sqrt[3]{(-3)^3} \implies -3\] do you get it?

OpenStudy (jiteshmeghwal9):

cube root of 27=3 cube root of -1=i

OpenStudy (lgbasallote):

no @jiteshmeghwal9

OpenStudy (jiteshmeghwal9):

3i=3*-1=-3

OpenStudy (jiteshmeghwal9):

that's ht i wanna say @lgbasallote

OpenStudy (lgbasallote):

cube root of -1 is -1 square root of -1 is i

OpenStudy (jiteshmeghwal9):

wht*

OpenStudy (jiteshmeghwal9):

oops sorry i don't see tht

OpenStudy (anonymous):

right when odd its the same sign but when even there opposite

OpenStudy (anonymous):

OK so WHY do you get -3 from the last equation you wrote.

OpenStudy (anonymous):

Oh wait it just dawned on me

OpenStudy (lgbasallote):

oh lol that's good

OpenStudy (anonymous):

its perfect square root thats why I was reading into something that was not there

OpenStudy (lgbasallote):

perfect cube root*

OpenStudy (anonymous):

oops duh lol

OpenStudy (lgbasallote):

hehe

OpenStudy (anonymous):

right you are somehow lucky to have such a large audience xD

OpenStudy (hba):

lol ^

OpenStudy (anonymous):

thanks @hba